# Symbiotic Game and Foundation Models for Cyber Deception Operations in Strategic Cyber Warfare

### v1

@@ Abstract
We are currently facing unprecedented cyber warfare with the rapid evolution of tactics, increasing asymmetry of intelligence, and the growing accessibility of hacking tools. In this landscape, cyber deception emerges as a critical component of our defense strategy against increasingly sophisticated attacks. This chapter aims to highlight the pivotal role of game-theoretic models and foundation models (FMs) in analyzing, designing, and implementing cyber deception tactics. Game models (GMs) serve as a foundational framework for modeling diverse adversarial interactions, allowing us to encapsulate both adversarial knowledge and domain-specific insights. Meanwhile, FMs serve as the building blocks for creating tailored machine learning models suited to given applications. By leveraging the synergy between GMs and FMs, we can advance proactive and automated cyber defense mechanisms by not only securing our networks against attacks but also enhancing their resilience against well-planned operations. This chapter discusses the games at the tactical, operational, and strategic levels of warfare, delves into the symbiotic relationship between these methodologies, and explores relevant applications where such a framework can make a substantial impact in cybersecurity. The chapter discusses the promising direction of the multi-agent neurosymbolic conjectural learning (MANSCOL), which allows the defender to predict adversarial behaviors, design adaptive defensive deception tactics, and synthesize knowledge for the operational level synthesis and adaptation. FMs serve as pivotal tools across various functions for MANSCOL, including reinforcement learning, knowledge assimilation, formation of conjectures, and contextual representation. This chapter concludes with the discussion of the challenges associated with the FMs and their application in the domain of cybersecurity.
@@ 1 Introduction
Cyber deception emerges as one promising class of defenses, involving the creation of traps, decoys, or false information to mislead attackers and divert their attention from valuable assets or genuine security measures [51, 1]. The deployment of cyber deception techniques enables network systems to detect, delay, or disrupt attackers’ activities while gathering intelligence about their tactics, techniques, and procedures (TTPs). An automated and adaptive cyber deception strategy can proactively predict and respond to attackers’ behaviors by creating further deception to deter the attack or engage attackers to gather information. By minimizing the involvement of human operators in the cyber response workflow and intelligently aligning tactical responses with mission objectives, this approach facilitates faster and more mission-aligned decision-making, enhancing cyber resilience for the mission.
@@ 1.1 Role of Game-Theoretic Models
The interaction between defenders and attackers in cyber deception and the cyber warfare in general is often conceptualized as a strategic game, wherein both parties endeavor to outmaneuver the other to achieve their respective goals. In recent literature, many game models (GMs) serve a descriptive framework, wherein a specific attack model, including strategies, objectives, and information structures, is assumed alongside a defender model [63]. Subsequently, the analysis of the game aims to find the equilibria of the game, which include equilibrium payoffs, strategies, and, in dynamic games, equilibrium dynamics or trajectories.
@@
This methodological framework finds applications across diverse domains, including jamming games [104, 106, 91, 71], authentication games [30, 82, 29, 34, 84], routing games [101, 109, 21], intrusion detection [80, 40, 100, 107], infrastructure protection games [16, 42, 12, 11, 41], and insider threat [47, 8, 9, 25, 65]. These endeavors yield valuable insights and establish a fundamental understanding of adversarial interactions at a higher level. Typically, the equilibrium analysis of these models plays a crucial role in informing strategic-level decisions, such as risk assessment, investment decisions, and resource planning within organizations and security contexts. Several factors contribute to the significance of equilibrium. Firstly, the credibility of equilibrium forecasts increases as games are played repeatedly over an extended period. With time, players tend to figure out the structure and the rule of the game, adopt rational strategies, and make more informed decision-making. Secondly, strategic-level decisions do not mandate a high-level precision of the details of implementation and operations and thus can rely on the equilibrium prediction. A rough estimate of the equilibrium outcomes within the correct magnitude suffices for effective high-level decision-making processes. One area for which game-theoretic frameworks have offered valuable tools is the strategic risk analysis against a variety of adversarial models, each representing contingent goal-driven adversaries [81, 19, 14]. This application is particularly vital within cybersecurity contexts, where predicting adversarial behaviors proves challenging, unlike natural events governed by probabilistic distributions. Adversarial TTPs drive these behaviors. Leveraging game theory provides a distinctive framework for cyber risk analysis, with implications spanning domains like cyber insurance [95, 93, 85, 94], cyber-physical system integration [46, 18, 108, 46], and security investment [10, 33, 15, 17].
@@
However, the lack of precision makes it less applicable to tactical scenarios where details of how to act are imperative in the face of an attack. Obtaining exact models of the game proves challenging, particularly regarding the strategies of attackers and their knowledge. Sometimes, it is also challenging for the defender itself to figure out the knowledge and available strategies. To address this challenge, literature often turns to more intricate modeling approaches, such as Bayesian games [36, 37] and hypergames [4, 88], to accommodate uncertainties. In particular, in the domain of cyber deception, where information asymmetry is prevalent, Bayesian games have emerged as an appropriate class for describing cyber deception scenarios. These scenarios involve various tactics such as honeypot configuration [54, 55, 5], attack engagement [45, 46, 7], and information design [98, 97, 96], among others.
@@
Moreover, while this approach is reasonable for handling structured uncertainties (i.e., known unknowns), it encounters difficulties with unstructured uncertainties (e.g., epistemic uncertainties). Furthermore, even with a well-defined model, computing equilibrium concepts like Bayesian equilibrium proves challenging [64] and often lacks predictive power, especially in dynamic tactical environments characterized by nonstationary behaviors and conditions.
@@
Learning in a nonstationary environment poses challenges as strategies adapted to earlier games must be re-evaluated for new environments [63]. Unlike in stationary environments, where learning strategies often converge to relevant equilibrium concepts [62], nonstationary games cannot be characterized by equilibrium, and there is a need for a reevaluation of reinforcement learning goals. Learning strategies do not need to aim to converge to an equilibrium outcome [66]; rather, an equilibrium outcome can be viewed as a result of learning behaviors [75]. This perspective can lead to the nonequilibrium solutions that redefine outcomes for nonstationary games [35]. Nonequilibrium solutions are relevant even in stationary games where interactions are limited, meaning players cannot reach equilibrium within short timeframes.
@@ 1.3 Cyber Deception and Related Game-Theoretic and Foundation Models
The essence of cyber deception resonates strongly with game-theoretic models. Cyber deception has essential characteristics that align closely with game-theoretic attributes, including the information asymmetry between the players, multi-stage and multi-phase interactions, the presence of aleatoric and epistemic uncertainties, together with the bounded rational human behaviors. In [77], a taxonomy of cyber deception games is developed to connect each deception scenario with a fundamental class of game-theoretic models. The taxonomy leads to a library of game building blocks that can be used to synthesize multi-round multi-scale dynamic games that capture the cyber kill and defense chains for strategic, operational, and tactic level decision-making. For example, honeypot engagement has been captured by a class of stochastic games with one-sided information. Honeypot deployment problems have been captured by network design games. Intrusion evasion games have been captured by a class of principal-agent detection games. These building blocks enable a divide-and-conquer and modular approach to design cyber defense. These modular games can be fused together with FMs to create function modules that are capable of representing games in the security scenario to provide strengthened learning and computation power. This enables the development of tactical-level applications and the enhancement of cyber deception techniques.
@@ 2 Cyber Deception and Network Security Games
Cyber deception stands out due to its unique tactics and objectives, which involve the strategic deployment of decoys, honeypots, and false information to manipulate attackers’ perceptions of the network and influence their actions. The primary aim of deception is twofold: to deter attacks from exploiting unknown vulnerabilities within our systems and to gain valuable intelligence from attackers’ behaviors. Game-theoretic models for cyber deception often revolve around information structures. For instance, asymmetric information games elucidate the information asymmetry between two players. Signaling games serve as a fundamental framework for understanding two-player deception scenarios, where one player observes the true state of the world while the other player can only observe messages or actions. They have been applied to study detection evasion [40], honeypots [76], and insider threats [8].
Stochastic games with one-sided information, e.g., in [39, 58, 2], extend this framework into dynamic settings, where one player has complete observation of the state space while the other player navigates with incomplete information, forming beliefs over the state space. Information acquisition games facilitate the trade-off between the cost of gathering information and the quality of decision-making based on such information. Recent literature has been investigating the design of observation kernels [67, 99] and optimal timing for information acquisition [49, 50, 48].
@@ 2.1 Multi-Level and Multi-Scale Game Models
As illustrated in Fig. 1, the GMs involved at the three levels are multi-scale. At the strategic level, the game is a high-level, coarse-grain description of the adversarial interactions between a defender and an attacker. For example, Blotto-type games, e.g., [83, 38, 31] have been classically used to allocate a limited amount of resources to overpower adversaries in cyber warfare by outnumbering their resources. FlipIt games and their variants have also been recently studied in [86, 56, 6] to strategically analyze the interplay between attackers and defenders, both striving to assert control over a resource for the maximum duration while minimizing overall move costs. Through strategic reasoning, defenders devise counterstrategies aimed at controlling resources at opportune moments and minimizing downtime and potential compromise. Network games are another class of strategic-level security games. Networks are used to capture the interdependencies and connectivity among the resources. The goal of the defender is to protect or defend essential network assets and minimize the impact of the attacks. For example, the framework proposed by [13] introduces a network design paradigm allowing network designers to fortify links to establish redundant communication pathways between nodes to mitigate potential attacks through resource investment. By proactively considering strategic cyber threats, the framework offers methods to characterize and compute optimal strategies for securing a network within predefined budget constraints. In addition to protective measures, resilient strategies for network recovery following attacks are also essential in network defense. A two-player, three-stage game framework proposed in [11] has aimed to capture both protection and recovery phases. The network designer’s objective is to maintain network connectivity both before and after an attack, while the adversary seeks to disrupt the network by compromising specific links. This strategic game framework facilitates analysis of trade-offs not only among nodes and links within the network but also among resources allocated for pre-event protection and post-event recovery.
@@
An example of the operation-level games is described in [105]. A multi-phase and multi-stage game is used to capture the attack vectors of advanced persistent threats. Each phase and each stage has distinct characteristics and requires a different form of defense strategy. A dynamic game framework can piece together games at each phase and each stage to capture the whole attack-and-defense operation holistically from the entry point to the target. A more sophisticated dynamic game that incorporates incomplete information and cyber deception has been presented in [46] and [45]. It depicts long-term dyadic interaction between a stealthy attacker and a proactive defender through a multi-stage game of incomplete information. Each player holds private information unknown to the other, and strategic actions are guided by beliefs formed through multi-stage observation and learning processes.
@@
The game at the tactic level is represented by the short-term local interactions between an attacker and a defender. Each stage of the operation involves a distinct type of dyadic interactions. For example, in the social engineering game at the reconnaissance stage, the attacker can exploit the cognitive vulnerabilities of the defenders to gain initial access. The attacker often has an information advantage when it aims to deceive the defender. At the stage of lateral movement, the defender can guide the attacker to a honeypot so that the attack will be detected and revealed. This defensive deception provides the defender information advantage at the stage of the game. Hence, throughout the operation, the game can take different forms depending on the structure of information, interactions, and maneuverability. The survey in [77] explores game theory frameworks designed to model defensive deception strategies in cybersecurity and privacy domains. It introduces a taxonomy comprising six distinct types of deception: perturbation, moving target defense, obfuscation, mixing, honey-x, and attacker engagement. These classifications are characterized by their information structures, agents involved, actions taken, and duration, all of which align closely with game theory concepts. This taxonomy establishes a structured framework for comprehending the diverse strategies of defensive deception. The survey outlined in [68] offers a curated selection of studies that demonstrate the application of game theory in addressing various security and privacy challenges within computer networks and mobile applications. In each domain, we identify security problems, players, and GMs, including security of the physical and MAC layers, security of self-organizing networks, intrusion detection systems, anonymity and privacy, economics of network security, and cryptography. The games that describe the local interaction can be viewed as a building block to synthesize an operational level game, and hence informing the creation of the strategic level games.
@@
Accompanied by the ability to zoom in and out, new solution concepts beyond equilibrium are necessary to predict game outcomes across various levels and resolutions. These concepts must be compatible with uncertainties, limited observations, and the non-stationary nature of interactions. Hence, nonequilibrium solutions are essential to address these features, aiming to create a reasonable prediction even though the outcome of the game can be stochastic with limited rounds of interactions. Moreover, solutions for tactical-level games must be consistent with those at the strategic level, despite differences in scales or levels of detail. This consistency represents another dimension of consistency, in addition to conjectural and incentive consistencies (e.g., associated with the solution concepts Berk-Nash equilibrium [24], Nash equilibrium [70]), leading to a holonic solution where localized tactical games consistent with global strategic ones.

### v2

@@ Abstract
We are currently facing unprecedented cyber warfare with the rapid evolution of tactics, increasing asymmetry of intelligence, and the growing accessibility of hacking tools. In this landscape, cyber deception emerges as a critical component of our defense strategy against increasingly sophisticated attacks. This chapter aims to highlight the pivotal role of game-theoretic models and foundation models (FMs) in analyzing, designing, and implementing cyber deception tactics. Game models (GMs) serve as a foundational framework for modeling diverse adversarial interactions, allowing us to encapsulate both adversarial knowledge and domain-specific insights. Meanwhile, FMs serve as the building blocks for creating tailored machine learning models suited to given applications. By leveraging the synergy between GMs and FMs, we can advance proactive and automated cyber defense mechanisms by not only securing our networks against attacks but also enhancing their resilience against well-planned operations. This chapter discusses the games at the tactical, operational, and strategic levels of warfare, delves into the symbiotic relationship between these methodologies, and explores relevant applications where such a framework can make a substantial impact in cybersecurity. The chapter discusses the promising direction of multi-agent neurosymbolic conjectural learning (MANSCOL), which allows the defender to predict adversarial behaviors, design adaptive defensive deception tactics, and synthesize knowledge for the operational level synthesis and adaptation. FMs serve as pivotal tools across various functions for MANSCOL, including reinforcement learning, knowledge assimilation, formation of conjectures, and contextual representation. This chapter concludes with a discussion of the challenges associated with FMs and their application in the domain of cybersecurity.
@@ 1 Introduction
Cyber deception emerges as one promising class of defenses, involving the creation of traps, decoys, or false information to mislead attackers and divert their attention from valuable assets or genuine security measures Jajodia et al. (2016); Al-Shaer et al. (2019). The deployment of cyber deception techniques enables network systems to detect, delay, or disrupt attackers’ activities while gathering intelligence about their tactics, techniques, and procedures (TTPs). An automated and adaptive cyber deception strategy can proactively predict and respond to attackers’ behaviors by creating further deception to deter the attack or engage attackers to gather information. By minimizing the involvement of human operators in the cyber response workflow and intelligently aligning tactical responses with mission objectives, this approach facilitates faster and more mission-aligned decision-making, enhancing cyber resilience for the mission.
@@ 1.1 Role of Game-Theoretic Models
The interaction between defenders and attackers in cyber deception and the cyber warfare in general is often conceptualized as a strategic game, wherein both parties endeavor to outmaneuver the other to achieve their respective goals. In recent literature, many game models (GMs) serve a descriptive framework, wherein a specific attack model, including strategies, objectives, and information structures, is assumed alongside a defender model Li et al. (2022). Subsequently, the analysis of the game aims to find the equilibria of the game, which include equilibrium payoffs, strategies, and, in dynamic games, equilibrium dynamics or trajectories.
@@
This methodological framework finds applications across diverse domains, including jamming games Zhu et al. (2010, 2011); Xu and Zhu (2017); Nugraha et al. (2019), authentication games Ge and Zhu (2023); Rass et al. (2020); Ge and Zhu (2022); Gupta et al. (2022); Sarıtaş et al. (2019), routing games Zhu et al. (2012a, b); Clark et al. (2012), intrusion detection Rass et al. (2017); Hu and Zhu (2022); Zhu and Başar (2009); Zhu et al. (2010), infrastructure protection games Chen and Zhu (2019); Huang et al. (2018); Chen et al. (2019, 2017); Huang et al. (2017), and insider threat Huang and Zhu (2021); Casey et al. (2016, 2015); Feng et al. (2015); Liu et al. (2008). These endeavors yield valuable insights and establish a fundamental understanding of adversarial interactions at a higher level. Typically, the equilibrium analysis of these models plays a crucial role in informing strategic-level decisions, such as risk assessment, investment decisions, and resource planning within organizations and security contexts. Several factors contribute to the significance of equilibrium. Firstly, the credibility of equilibrium forecasts increases as games are played repeatedly over an extended period. With time, players tend to figure out the structure and the rule of the game, adopt rational strategies, and make more informed decision-making. Secondly, strategic-level decisions do not mandate a high-level precision of the details of implementation and operations and thus can rely on the equilibrium prediction. A rough estimate of the equilibrium outcomes within the correct magnitude suffices for effective high-level decision-making processes. One area for which game-theoretic frameworks have offered valuable tools is the strategic risk analysis against a variety of adversarial models, each representing contingent goal-driven adversaries Rass and Schauer (2018); Chen et al. (2021); Chen and Zhu (2018). This application is particularly vital within cybersecurity contexts, where predicting adversarial behaviors proves challenging, unlike natural events governed by probabilistic distributions. Adversarial TTPs drive these behaviors. Leveraging game theory provides a distinctive framework for cyber risk analysis, with implications spanning domains like cyber insurance Zhang et al. (2017); Zhang and Zhu (2019); Schwartz and Sastry (2014); Zhang and Zhu (2021), cyber-physical system integration Huang and Zhu (2020); Chen and Zhu (2022); Zhu and Xu (2020); Huang and Zhu (2020), and security investment Cavusoglu et al. (2008); Grossklags and Johnson (2009); Chen and Zhu (2018, 2019).
@@
However, the lack of precision makes it less applicable to tactical scenarios where details of how to act are imperative in the face of an attack. Obtaining exact models of the game proves challenging, particularly regarding the strategies of attackers and their knowledge. Sometimes, it is also challenging for the defender itself to figure out the knowledge and available strategies. To address this challenge, literature often turns to more intricate modeling approaches, such as Bayesian games Harsanyi (1968, 1995) and hypergames Bennett (1980); Wang et al. (1989), to accommodate uncertainties. In particular, in the domain of cyber deception, where information asymmetry is prevalent, Bayesian games have emerged as an appropriate class for describing cyber deception scenarios. These scenarios involve various tactics such as honeypot configuration La et al. (2016a, b); Boumkheld et al. (2019), attack engagement Huang and Zhu (2019, 2020); Caballero et al. (2024), and information design Zhang and Zhu (2023, 2022, 2021), among others.
@@
Moreover, while this approach is reasonable for handling structured uncertainties (i.e., known unknowns), it encounters difficulties with unstructured uncertainties (e.g., epistemic uncertainties). Furthermore, even with a well-defined model, computing equilibrium concepts like Bayesian equilibrium proves challenging Li and Zhu (2023) and often lacks predictive power, especially in dynamic tactical environments characterized by nonstationary behaviors and conditions.
@@
Learning in a nonstationary environment poses challenges as strategies adapted to earlier games must be re-evaluated for new environments Li et al. (2022). Unlike in stationary environments, where learning strategies often converge to relevant equilibrium concepts Li et al. (2022), nonstationary games cannot be characterized by equilibrium, and there is a need for a reevaluation of reinforcement learning goals. Learning strategies do not need to aim to converge to an equilibrium outcome Liu et al. (2023); rather, an equilibrium outcome can be viewed as a result of learning behaviors Pan et al. (2023). This perspective can lead to the nonequilibrium solutions that redefine outcomes for nonstationary games Hammar et al. (2024). Nonequilibrium solutions are relevant even in stationary games where interactions are limited, meaning players cannot reach equilibrium within short timeframes.
@@ 1.3 Cyber Deception and Related Game-Theoretic and Foundation Models
The essence of cyber deception resonates strongly with game-theoretic models. Cyber deception has essential characteristics that align closely with game-theoretic attributes, including the information asymmetry between the players, multi-stage and multi-phase interactions, the presence of aleatoric and epistemic uncertainties, together with the bounded rational human behaviors. In Pawlick et al. (2019), a taxonomy of cyber deception games is developed to connect each deception scenario with a fundamental class of game-theoretic models. The taxonomy leads to a library of game building blocks that can be used to synthesize multi-round multi-scale dynamic games that capture the cyber kill and defense chains for strategic, operational, and tactic level decision-making. For example, honeypot engagement has been captured by a class of stochastic games with one-sided information. Honeypot deployment problems have been captured by network design games. Intrusion evasion games have been captured by a class of principal-agent detection games. These building blocks enable a divide-and-conquer and modular approach to design cyber defense. These modular games can be fused together with FMs to create function modules that are capable of representing games in the security scenario to provide strengthened learning and computation power. This enables the development of tactical-level applications and the enhancement of cyber deception techniques.
@@ 2 Cyber Deception and Network Security Games
Cyber deception stands out due to its unique tactics and objectives, which involve the strategic deployment of decoys, honeypots, and false information to manipulate attackers’ perceptions of the network and influence their actions. The primary aim of deception is twofold: to deter attacks from exploiting unknown vulnerabilities within our systems and to gain valuable intelligence from attackers’ behaviors. Game-theoretic models for cyber deception often revolve around information structures. For instance, asymmetric information games elucidate the information asymmetry between two players. Signaling games serve as a fundamental framework for understanding two-player deception scenarios, where one player observes the true state of the world while the other player can only observe messages or actions. They have been applied to study detection evasion Hu and Zhu (2022), honeypots Pawlick et al. (2018), and insider threats Casey et al. (2016).
Stochastic games with one-sided information, e.g., in Horák et al. (2017); Li et al. (2024); Anwar and Kamhoua (2020), extend this framework into dynamic settings, where one player has complete observation of the state space while the other player navigates with incomplete information, forming beliefs over the state space. Information acquisition games facilitate the trade-off between the cost of gathering information and the quality of decision-making based on such information. Recent literature has been investigating the design of observation kernels Liu and Zhu (2023); Zhang and Zhu (2020) and optimal timing for information acquisition Huang et al. (2019); Huang and Zhu (2021); Huang et al. (2021).
@@ 2.1 Multi-Level and Multi-Scale Game Models
As illustrated in Fig. 1, the GMs involved at the three levels are multi-scale. At the strategic level, the game is a high-level, coarse-grain description of the adversarial interactions between a defender and an attacker. For example, Blotto-type games, e.g., Roberson (2006); Hart (2008); Golman and Page (2009) have been classically used to allocate a limited amount of resources to overpower adversaries in cyber warfare by outnumbering their resources. FlipIt games and their variants have also been recently studied in Van Dijk et al. (2013); Laszka et al. (2014); Bowers et al. (2012) to strategically analyze the interplay between attackers and defenders, both striving to assert control over a resource for the maximum duration while minimizing overall move costs. Through strategic reasoning, defenders devise counterstrategies aimed at controlling resources at opportune moments and minimizing downtime and potential compromise. Network games are another class of strategic-level security games. Networks are used to capture the interdependencies and connectivity among the resources. The goal of the defender is to protect or defend essential network assets and minimize the impact of the attacks. For example, the framework proposed by Chen et al. (2020) introduces a network design paradigm allowing network designers to fortify links to establish redundant communication pathways between nodes to mitigate potential attacks through resource investment. By proactively considering strategic cyber threats, the framework offers methods to characterize and compute optimal strategies for securing a network within predefined budget constraints. In addition to protective measures, resilient strategies for network recovery following attacks are also essential in network defense. A two-player, three-stage game framework proposed in Chen et al. (2017) has aimed to capture both protection and recovery phases. The network designer’s objective is to maintain network connectivity both before and after an attack, while the adversary seeks to disrupt the network by compromising specific links. This strategic game framework facilitates analysis of trade-offs not only among nodes and links within the network but also among resources allocated for pre-event protection and post-event recovery.
@@
An example of the operation-level games is described in Zhu and Rass (2018). A multi-phase and multi-stage game is used to capture the attack vectors of advanced persistent threats. Each phase and each stage has distinct characteristics and requires a different form of defense strategy. A dynamic game framework can piece together games at each phase and each stage to capture the whole attack-and-defense operation holistically from the entry point to the target. A more sophisticated dynamic game that incorporates incomplete information and cyber deception has been presented in Huang and Zhu (2020) and Huang and Zhu (2019). It depicts long-term dyadic interaction between a stealthy attacker and a proactive defender through a multi-stage game of incomplete information. Each player holds private information unknown to the other, and strategic actions are guided by beliefs formed through multi-stage observation and learning processes.
@@
The game at the tactic level is represented by the short-term local interactions between an attacker and a defender. Each stage of the operation involves a distinct type of dyadic interactions. For example, in the social engineering game at the reconnaissance stage, the attacker can exploit the cognitive vulnerabilities of the defenders to gain initial access. The attacker often has an information advantage when it aims to deceive the defender. At the stage of lateral movement, the defender can guide the attacker to a honeypot so that the attack will be detected and revealed. This defensive deception provides the defender information advantage at the stage of the game. Hence, throughout the operation, the game can take different forms depending on the structure of information, interactions, and maneuverability. The survey in Pawlick et al. (2019) explores game theory frameworks designed to model defensive deception strategies in cybersecurity and privacy domains. It introduces a taxonomy comprising six distinct types of deception: perturbation, moving target defense, obfuscation, mixing, honey-x, and attacker engagement. These classifications are characterized by their information structures, agents involved, actions taken, and duration, all of which align closely with game theory concepts. This taxonomy establishes a structured framework for comprehending the diverse strategies of defensive deception. The survey outlined in Manshaei et al. (2013) offers a curated selection of studies that demonstrate the application of game theory in addressing various security and privacy challenges within computer networks and mobile applications. In each domain, we identify security problems, players, and GMs, including security of the physical and MAC layers, security of self-organizing networks, intrusion detection systems, anonymity and privacy, economics of network security, and cryptography. The games that describe the local interaction can be viewed as a building block to synthesize an operational level game, and hence informing the creation of the strategic level games.
@@
Accompanied by the ability to zoom in and out, new solution concepts beyond equilibrium are necessary to predict game outcomes across various levels and resolutions. These concepts must be compatible with uncertainties, limited observations, and the non-stationary nature of interactions. Hence, nonequilibrium solutions are essential to address these features, aiming to create a reasonable prediction even though the outcome of the game can be stochastic with limited rounds of interactions. Moreover, solutions for tactical-level games must be consistent with those at the strategic level, despite differences in scales or levels of detail. This consistency represents another dimension of consistency, in addition to conjectural and incentive consistencies (e.g., associated with the solution concepts Berk-Nash equilibrium Esponda and Pouzo (2016), Nash equilibrium Nash (1953)), leading to a holonic solution where localized tactical games consistent with global strategic ones.